Conservation of Energy Ranking Task Six pendulums of various masses m are released from various heights h above a tabletop, as shown in the figures below. All the pendulums have the same length and are mounted such that at the vertical position their lowest points are the height of the tabletop and just do not strike the tabletop when released. Assume that the size of each bob is negligible.

Rank each pendulum on the basis of its initial gravitational potential energy (before being released) relative to the tabletop.

Rank from largest to smallest To rank items as equivalent, overlap them.
Answer #1
Concepts and reason

The concept used to solve this problem is gravitational potential energy.

Find the gravitational potential energy in each of the six cases and arrange in the increasing order of gravitational potential energy.

Fundamentals

The gravitational potential energy is the energy that an object possesses due to its position above the surface of the earth.

The equation for gravitational potential energy is,

P= mgh

Here gravitational potential energy is , mass is , acceleration due to gravity is and height is .

The gravitational potential energy is,

P= mgh

Substituting for , 9.8m/s2
for and 60 cm
for .

(9LTIN
(0)(u009)(_srug:o)(3x7)=d

The gravitational potential energy is given by the expression

Substituting for , 9.8m/s2
for and for .

P=(1kg)(9.8 m/s?)(60cm) 100cm
100 cm
= 5.88J

The gravitational potential energy is given by the expression

Substituting for , 9.8m/s2
for and for .

P=(4kg)(9.8m/s?)(30cm) 100cm)
= 11.76J

The gravitational potential energy is given by the expression

Substituting for , 9.8m/s2
for and for .

P=(2x3)(9.8m?)(30cm) 100mm
= 5.88J

The gravitational potential energy is given by the expression

Substituting for , 9.8m/s2
for and for .

( lm
P=(8kg)(9.8 m/s*)(15cm)
1812.0m (100cm
= 11.76J

The gravitational potential energy is given by the expression

Substituting for , 9.8m/s2
for and 45cm
for .

(Im
P=(3kg)(9.8m/s?
KB)(9.8 m/s)(45cm) (100cm)
= 13.23J

Ans:

The gravitational potential energy can be arranged in the increasing order as follows:

P(m = 3kg, h = 45cm) > P(m = 2 kg, h = 60cm)=P(m=4kg, h = 30cm)=P(m=8kg, h=15cm)
> P(m=1kg, h = 60cm)=P(m = 2 kg, h = 30cm)